# library.bib

### v1

@@ A.3 Additional Experimental Results
Table 11: Statistics for predicting $\vec{v}$ with the randomK dataset. Errors in [m/y], MSE in [m2/y2], SSIM unitless. Mean $\pm$ standard deviation reported.
Model Data Output Case Huber L MAE MSE SSIM
LGCNN randomK3 $v_{x}$ train 14.11 $\pm$ 11.52 96.63 $\pm$ 23.29 14.60 $\pm$ 11.53 380.19 $\pm$ 529.31 0.997 $\pm$ 0.002
$v_{y}$ train 14.30 $\pm$ 7.10 135.26 $\pm$ 20.36 14.80 $\pm$ 7.11 361.61 $\pm$ 296.21 0.991 $\pm$ 0.006
$v_{x}$ val 28.23 $\pm$ 7.76 445.13 $\pm$ 35.50 28.72 $\pm$ 7.77 1601.10 $\pm$ 587.80 0.989 $\pm$ 0.003
$v_{y}$ val 27.33 $\pm$ 2.25 343.35 $\pm$ 26.80 27.82 $\pm$ 2.25 1587.57 $\pm$ 252.10 0.982 $\pm$ 0.002
$v_{x}$ test 26.94 $\pm$ 5.99 216.74 $\pm$ 18.92 27.44 $\pm$ 5.99 1368.52 $\pm$ 431.48 0.990 $\pm$ 0.002
$v_{y}$ test 30.38 $\pm$ 6.18 249.11 $\pm$ 11.37 30.88 $\pm$ 6.18 1848.47 $\pm$ 637.10 0.982 $\pm$ 0.005
$v_{x}$ scaling 27.89 $\pm$ 6.53 286.08 $\pm$ 22.96 28.38 $\pm$ 6.53 1486.23 $\pm$ 451.52 0.990 $\pm$ 0.002
$v_{y}$ scaling 27.60 $\pm$ 2.43 328.80 $\pm$ 24.95 28.09 $\pm$ 2.43 1578.87 $\pm$ 258.66 0.979 $\pm$ 0.003
LGCNN realK $v_{x}$ train 29.66 $\pm$ 2.29 592.17 $\pm$ 37.66 30.16 $\pm$ 2.29 2326.33 $\pm$ 327.11 0.988 $\pm$ 0.005
$v_{y}$ train 28.93 $\pm$ 5.47 397.09 $\pm$ 72.78 29.43 $\pm$ 5.47 1778.28 $\pm$ 551.97 0.997 $\pm$ 0.009
$v_{x}$ val 22.17 $\pm$ 2.24 170.47 $\pm$ 9.75 22.66 $\pm$ 2.24 909.39 $\pm$ 139.94 0.988 $\pm$ 0.002
$v_{y}$ val 20.20 $\pm$ 4.07 113.33 $\pm$ 12.37 20.70 $\pm$ 4.08 645.11 $\pm$ 221.50 0.997 $\pm$ 0.001
$v_{x}$ scaling 79.33 $\pm$ 21.80 532.35 $\pm$ 140.50 79.83 $\pm$ 21.80 9404.62 $\pm$ 4048.09 0.953 $\pm$ 0.014
$v_{y}$ scaling 53.39 $\pm$ 12.64 602.41 $\pm$ 128.79 53.89 $\pm$ 12.64 4621.25 $\pm$ 1755.50 0.980 $\pm$ 0.006
@@
The results of the LGCNN and DDU-Net, evaluated on both synthetic and realistic permeability fields, for the training, validation, testing, and scaling datasets are shown in Table 12 (for Step 1 - predict $\vec{v}_{\text{pred}}$) and Table 13 (for Step 3 - predict $T$). The column "data" refers to the dataset that a model was trained on and applied to, i.e., "randomK" stands for the baseline dataset 3dp, "randomK101" for 101dp. An asterisk indicates that the model was evaluated on a dataset different from the one it was trained on. Mostly relevant for the vanilla approaches that are trained on 101dp and applied to scaling and test of 3dp.
@@
Table 12: Performance metrics for for predicting $\vec{v}$ with different models and datasets. Errors in [m/y], MSE in [m2/y2], SSIM unitless.
Model Data Output Case Huber L MAE MSE SSIM
1st Step
LGCNN randomK $v_{x}$ train 9.4732 132.6672 9.9620 171.0032 0.9972
$v_{y}$ train 11.4005 223.4601 11.8902 275.6190 0.9937
$v_{x}$ val 22.2241 343.9907 22.7179 1102.3721 0.9905
$v_{y}$ val 26.6078 274.8036 27.1026 1524.5099 0.9841
$v_{x}$ test 21.8237 190.8046 22.3178 972.5668 0.9911
$v_{y}$ test 32.2488 256.2519 32.7444 2031.3357 0.9812
* $v_{x}$ scaling 24.4314 294.0457 24.9261 1204.1154 0.9911
* $v_{y}$ scaling 25.7847 367.6891 26.2795 1463.8218 0.9820
Experiment: trained on full image
LGCNN randomK $v_{x}$ train 3.1983 39.9380 3.6714 21.5058 0.9993
$v_{y}$ train 3.0496 43.9018 3.5177 21.4905 0.9988
$v_{x}$ val 25.2418 417.1447 25.7367 1421.6547 0.9896
$v_{y}$ val 29.6294 276.8029 30.1247 1827.3134 0.9824
$v_{x}$ test 23.3646 202.6397 23.8584 1144.4479 0.9913
$v_{y}$ test 35.4289 332.9557 35.9249 2490.4104 0.9789
Replace with UNet-101dp, i.e., $1\times 1$ subdomain-DDUNet, trained on 101 datapoints
DDUNet randomK101 $v_{x}$ train 3.5468 678.7761 4.0134 86.6731 0.9767
$v_{y}$ train 4.1030 690.1506 4.5816 100.2767 0.9934
$v_{x}$ val 3.8368 647.1833 4.3058 90.3292 0.9739
$v_{y}$ val 4.3649 685.6656 4.8433 102.8601 0.9931
$v_{x}$ test 3.6643 584.8119 4.1325 83.4062 0.9731
$v_{y}$ test 4.1355 547.8865 4.6131 87.7819 0.9924
* $v_{x}$ LGCNN-test 12.2161 1047.8464 12.7036 896.2758 0.9016
* $v_{y}$ LGCNN-test 13.9033 2594.9504 14.3878 2317.6863 0.9903
* $v_{x}$ scaling 4.7412 659.8439 5.2152 79.2508 0.9699
* $v_{y}$ scaling 5.2686 456.1904 5.7492 83.9735 0.9948
Replace with 2x2-DDUNet-101dp, i.e.,2×$\times 2$ 2 subdomains
DDUNet randomK101 $v_{x}$ train 5.3928 422.4526 5.8815 57.6045 0.9580
$v_{y}$ train 4.0476 696.2919 4.5116 58.8457 0.9972
$v_{x}$ val 5.7663 471.6309 6.2545 76.4372 0.9552
$v_{y}$ val 4.5085 644.1844 4.9751 71.5679 0.9973
$v_{x}$ test 5.5847 319.9844 6.0735 59.2501 0.9526
$v_{y}$ test 4.2292 535.0656 4.6948 54.5020 0.9972
* $v_{x}$ LGCNN-test 13.1829 2453.5881 13.6735 1367.9061 0.9325
* $v_{y}$ LGCNN-test 14.6481 2029.0269 15.1332 2221.4226 0.9919
* $v_{x}$ scaling 7.6311 421.4186 8.1184 125.9511 0.9314
* $v_{y}$ scaling 6.9406 400.2490 7.4200 135.7929 0.9955
Domain Transfer, i.e., to more realistic data
LGCNN realK $v_{x}$ train 13.6890 122.7981 14.1819 340.6713 0.9973
$v_{y}$ train 9.0680 71.2576 9.5619 126.7222 0.9991
$v_{x}$ val 14.9187 106.8607 15.4095 380.3762 0.9939
$v_{y}$ val 10.1675 74.5570 10.6605 148.2475 0.9993
* $v_{x}$ scaling 109.5079 459.1620 110.0078 13570.0000 0.9462
* $v_{y}$ scaling 17.1118 240.9406 17.6051 616.5186 0.9965
Table 13: Performance metrics for predicting $T$ with different models and datasets. Errors in [°C], MSE in [°C2], PATs in [%], SSIM unitless.
Model Data Case Huber L MAE MSE PAT SSIM
3rd Step: trained and applied to $v_{sim}$
LGCNN randomK $v_{sim}$ train 1.87e-5 2.2536 0.0014 3.95e-5 0.03 0.9997
val 0.0027 2.8857 0.0369 0.0054 8.61 0.9283
test 0.0021 2.8990 0.0347 0.0041 7.54 0.9304
* scaling 0.0007 3.0250 0.0168 0.0014 2.05 0.9510
Full Pipeline, i.e., Step 3 trained on $v_{sim}$, but applied to $v_{pred}$
LGCNN randomK $v_{pred}$ train 0.0121 5.1006 0.0642 0.0272 18.17 0.8700
val 0.0188 4.2264 0.0967 0.0411 28.96 0.7625
test 0.0147 4.2120 0.0905 0.0307 28.92 0.7637
* scaling 0.0065 4.9366 0.0413 0.0141 10.87 0.8654
Experiments on 3rd Step
Trained in sequence, i.e., trained on $v_{pred}$
LGCNN randomK $v_{pred}$ train 0.0032 4.0750 0.0198 0.0068 3.73 0.9794
val 0.0195 3.8178 0.1003 0.0423 30.46 0.7442
test 0.0139 4.0804 0.0901 0.0289 29.44 0.7553
Trained with zero-padding
LGCNN randomK $v_{sim}$ train 0.0002 1.8465 0.0140 0.0004 0.33 0.9518
val 0.0032 2.9051 0.0465 0.0064 11.31 0.8686
test 0.0024 2.3391 0.0418 0.0048 9.19 0.8838
Trained on full image
LGCNN randomK $v_{sim}$ train 0.0002 3.5648 0.0022 0.0004 0.09 0.9992
val 0.0027 3.5088 0.0375 0.0055 8.81 0.9217
test 0.0021 3.6420 0.0351 0.0043 7.41 0.9245
Replace with UNet-101dp: Step 3 trained and applied on $v_{sim}$
DDUNet rK101 train 0.0019 3.8491 0.0283 0.0038 5.35 0.9911
val 0.0027 3.9381 0.0350 0.0055 8.89 0.9895
test 0.0025 3.7999 0.0343 0.0052 8.50 0.9903
* LGCNN-test 0.0032 4.4658 0.0416 0.0067 11.69 0.9835
* scaling 0.0010 4.3489 0.0176 0.0022 3.07 0.9954
Replace with 2x2-DDUNet-101dp: Step 3 trained and applied on $v_{sim}$
DDUNet rK101 train 0.0015 4.0624 0.0200 0.0032 3.44 0.9952
val 0.0040 4.1863 0.0410 0.0083 12.54 0.9904
test 0.0039 4.0807 0.0411 0.0081 12.67 0.9910
* LGCNN-test 0.0029 4.2772 0.0376 0.0058 10.96 0.9915
* scaling 0.0012 3.9946 0.0192 0.0025 4.73 0.9968
Domain Transfer, i.e. to more realistic data
3rd Step
LGCNN realK $v_{sim}$ train 0.0002 0.7704 0.0139 0.0005 0.43 2.9107
val 0.0005 0.8222 0.0175 0.0010 2.18 0.9672
* scaling 0.0004 0.8052 0.0189 0.0008 0.92 0.9497
Full Pipeline
LGCNN realK $v_{pred}$ train 0.0049 2.5437 0.0534 0.0100 17.58 2.4923
val 0.0137 2.3194 0.0841 0.0275 27.79 0.7510
* scaling 0.0022 2.0511 0.0394 0.0044 10.02 0.8708
@@
The model generally obtains good results in Figure 3, even for cells that are far away from injection points.
@@
Additional tests show that excluding other inputs, either $i$ alone or both $i$ and $k$, raises prediction error by 58–121%. We also evaluated alternative time-stepping schemes for solving the IVP. Replacing the 5th-order implicit Runge–Kutta method with explicit 2nd- or 4th-order schemes accelerates computation, but increases prediction error by 16–23%—a moderate degradation compared to the complete removal of streamline inputs. Nonetheless, we retain the implicit scheme for its superior accuracy and stability. Quantitative and qualitative results for the predictions are shown in Figure 7 and Table 14.
Table 14: Experiment on 3rd step: Test metrics for predicting $T$ with different input combinations. Errors in [°C], MSE in [°C2], PAT in [%], SSIM unitless.
Inputs Huber L MAE MSE PAT SSIM
$ikv_{x}v_{y}$ 0.0070 2.2990 0.0712 0.0139 25.55 0.7662
$ikv_{x}v_{y}s$ 0.0041 1.8674 0.0545 0.0083 20.56 0.8368
$ikv_{x}v_{y}s_{o}$ 0.0057 2.6623 0.0598 0.0114 20.64 0.8423
$ikv_{x}v_{y}ss_{o}$ (not faded) 0.0072 2.3039 0.0744 0.0144 21.40 0.7837
$ikv_{x}v_{y}ss_{o}^{\mathrm{a}}$ 0.0031 1.8364 0.0442 0.0062 12.97 0.8828
$v_{x}v_{y}ss_{o}$ 0.0066 2.1301 0.0647 0.0132 20.64 0.8681
$kv_{x}v_{y}ss_{o}$ 0.0049 2.0925 0.0587 0.0097 20.32 0.8732
explicit RK, order 4 0.0038 3.2636 0.0486 0.0076 15.80 0.8871
explicit RK, order 2 0.0036 2.2034 0.0463 0.0072 13.75 0.8830
@@
with Huber validation loss, hence the results differ slightly wrt. to Table 3.
Figure 7: 1st Column: Label, input $k$, 2nd-5th: Without $(s,s_{o})$, $(s_{o})$, $(s)$, not-faded streamlines $(s,s_{o})$, 6th: include all inputs, 7th-8th: Without $(i,k)$, $(i)$. Absolute errors capped at 1°C for better visualizations. Maximum errors are listed in Table 14.
@@
The qualitative performance is observable in Figure 8, where we see coherent streamlines and plume structures for both the isolated 3rd step and the full pipeline.
@@
Table 15 summarizes the training and inference times, number of epochs, and dataset splits (train:val:test) for each of the three steps in our pipeline, both for the LGCNN trained on partitioned and full datasets. Although both approaches exhibit similar inference times, they show significant differences in training time, with the partitioned approach yielding better results.
Table 15: Training measurements on the randomK dataset.
@@
In Table 16, the number of epochs and the total training time for the data-driven approaches are shown. For Step 1, both the UNet and DDUNet need many epochs and comparable training time to converge; however, for the third step and the full pipeline, the DDUNet significantly reduces both the number of epochs and the total training time required to reach convergence.
@@
2$\times$2 DDUNet101dp 735 4.787 hours
@@
2$\times$2 DDUNet101dp 303 3.508 hours
@@
2$\times$2 DDUNet101dp 97 1.222 hours

### v2

@@ A.3 Additional Experimental Results
Table 8: Statistical error metrics for predicting with different models and datasets. Errors in [°C] or [m/y], MoC in [%], and SSIM unitless. Mean $\pm$ standard deviation reported.
Architecture Data Split Scaling Step MAE RMSE Huber SSIM MoC
UNeta synth.: 3dp train No end-to-end 0.0155 $\pm$ 0.0064 0.0732 $\pm$ 0.0310 0.0025 $\pm$ 0.0021 0.9645 $\pm$ 0.0157 1.6770 $\pm$ 1.9093
UNeta synth.: 3dp val No end-to-end 0.1114 $\pm$ 0.0021 0.2177 $\pm$ 0.0039 0.0219 $\pm$ 0.0007 0.6251 $\pm$ 0.0067 6.9180 $\pm$ 0.7236
UNeta synth.: 3dp test No end-to-end 0.1043 $\pm$ 0.0016 0.1917 $\pm$ 0.0035 0.0176 $\pm$ 0.0006 0.6239 $\pm$ 0.0048 7.7775 $\pm$ 1.5033
LGCNNb synth.: 3dp train No end-to-end 0.0720 $\pm$ 0.0012 0.1736 $\pm$ 0.0019 0.0137 $\pm$ 0.0003 0.7993 $\pm$ 0.0044 0.0933 $\pm$ 0.0430
LGCNNb synth.: 3dp val No end-to-end 0.1026 $\pm$ 0.0014 0.2095 $\pm$ 0.0032 0.0204 $\pm$ 0.0006 0.6812 $\pm$ 0.0061 1.6610 $\pm$ 0.0797
LGCNNb synth.: 3dp test No end-to-end 0.0965 $\pm$ 0.0013 0.1866 $\pm$ 0.0030 0.0167 $\pm$ 0.0005 0.6724 $\pm$ 0.0059 0.1443 $\pm$ 0.0388
LGCNNb synth.: 3dp test Yes end-to-end 0.0430 $\pm$ 0.0007 0.1249 $\pm$ 0.0016 0.0072 $\pm$ 0.0002 0.7908 $\pm$ 0.0151 0.0951 $\pm$ 0.0317
UNet synth.: 3dp train No 3 0.0354 $\pm$ 0.0007 0.0770 $\pm$ 0.0024 0.0029 $\pm$ 0.0002 0.9084 $\pm$ 0.0032 0.1516 $\pm$ 0.0203
UNet synth.: 3dp val No 3 0.0491 $\pm$ 0.0005 0.0916 $\pm$ 0.0008 0.0042 $\pm$ 0.0001 0.8251 $\pm$ 0.0033 0.1970 $\pm$ 0.0381
UNet synth.: 3dp test No 3 0.0452 $\pm$ 0.0005 0.0829 $\pm$ 0.0011 0.0034 $\pm$ 0.0001 0.8381 $\pm$ 0.0040 0.1938 $\pm$ 0.0907
UNet synth.: 3dp test Yes 3 0.0179 $\pm$ 0.0005 0.0399 $\pm$ 0.0007 0.0008 $\pm$ 0.0000 0.8809 $\pm$ 0.0138 0.1148 $\pm$ 0.0497
UNet synth.: 3dp train No 1 (vx) 14.6004 $\pm$ 10.3108 16.7442 $\pm$ 9.9912 14.1096 $\pm$ 10.3058 0.9884 $\pm$ 0.0020 -
UNet synth.: 3dp train No 1 (vy) 14.7965 $\pm$ 6.3568 17.9164 $\pm$ 6.3728 14.3037 $\pm$ 6.3526 0.9868 $\pm$ 0.0051 -
UNet synth.: 3dp val No 1 (vx) 28.7237 $\pm$ 6.9461 39.5616 $\pm$ 5.9981 28.2283 $\pm$ 6.9444 0.9734 $\pm$ 0.0025 -
UNet synth.: 3dp val No 1 (vy) 27.8241 $\pm$ 2.0130 39.7497 $\pm$ 2.7446 27.3289 $\pm$ 2.0125 0.9743 $\pm$ 0.0020 -
UNet synth.: 3dp test No 1 (vx) 27.4355 $\pm$ 5.3603 36.6786 $\pm$ 4.8168 26.9402 $\pm$ 5.3588 0.9727 $\pm$ 0.0025 -
UNet synth.: 3dp test No 1 (vy) 30.8755 $\pm$ 5.5264 42.5114 $\pm$ 6.4231 30.3799 $\pm$ 5.5256 0.9734 $\pm$ 0.0042 -
UNet synth.: 3dp test Yes 1 (vx) 28.3820 $\pm$ 5.8382 38.2430 $\pm$ 4.8683 27.8864 $\pm$ 5.8367 0.9770 $\pm$ 0.0017 -
UNet synth.: 3dp test Yes 1 (vy) 28.0927 $\pm$ 2.1725 39.6320 $\pm$ 2.8592 27.5972 $\pm$ 2.1719 0.9717 $\pm$ 0.0025 -
LGCNNb real: 4dp train No end-to-end 0.0550 $\pm$ 0.0025 0.0960 $\pm$ 0.0014 0.0046 $\pm$ 0.0001 0.6744 $\pm$ 0.0249 1.8951 $\pm$ 0.6214
LGCNNb real: 4dp val No end-to-end 0.0827 $\pm$ 0.0029 0.1553 $\pm$ 0.0031 0.0120 $\pm$ 0.0005 0.6337 $\pm$ 0.0195 2.0679 $\pm$ 0.4008
LGCNNb real: 4dp test Yes end-to-end 0.0486 $\pm$ 0.0075 0.0714 $\pm$ 0.0056 0.0026 $\pm$ 0.0004 0.6257 $\pm$ 0.0396 2.7049 $\pm$ 0.6239
UNet real: 4dp train No 3 0.0213 $\pm$ 0.0051 0.0295 $\pm$ 0.0061 0.0005 $\pm$ 0.0002 0.8895 $\pm$ 0.0208 1.9966 $\pm$ 0.2618
UNet real: 4dp val No 3 0.0239 $\pm$ 0.0023 0.0395 $\pm$ 0.0034 0.0008 $\pm$ 0.0001 0.8982 $\pm$ 0.0169 2.9922 $\pm$ 0.3038
UNet real: 4dp test Yes 3 0.0218 $\pm$ 0.0042 0.0326 $\pm$ 0.0044 0.0005 $\pm$ 0.0001 0.8734 $\pm$ 0.0190 1.6682 $\pm$ 0.2054
UNet real: 4dp train No 1 (vx) 22.0019 $\pm$ 5.7393 27.7180 $\pm$ 6.8232 21.5061 $\pm$ 5.7380 0.9731 $\pm$ 0.0280 -
UNet real: 4dp train No 1 (vy) 13.4570 $\pm$ 9.2203 16.8320 $\pm$ 10.8123 12.9699 $\pm$ 9.2128 0.9607 $\pm$ 0.0629 -
UNet real: 4dp val No 1 (vx) 24.8187 $\pm$ 4.0315 29.6583 $\pm$ 4.7384 24.3217 $\pm$ 4.0310 0.9641 $\pm$ 0.0220 -
UNet real: 4dp val No 1 (vy) 12.9956 $\pm$ 9.9800 15.4509 $\pm$ 10.3446 12.5082 $\pm$ 9.9727 0.9485 $\pm$ 0.0748 -
UNet real: 4dp test Yes 1 (vx) 122.7812 $\pm$ 32.6445 135.7671 $\pm$ 36.8100 122.2815 $\pm$ 32.6443 0.9128 $\pm$ 0.0511 -
UNet real: 4dp test Yes 1 (vy) 25.5754 $\pm$ 6.2873 34.6947 $\pm$ 8.2358 25.0801 $\pm$ 6.2858 0.9445 $\pm$ 0.0834 -
a newly trained with other validation loss for comparability in this table
b with fixed first step predictions
@@
The results of all steps/end-to-end, evaluated on both synthetic and realistic permeability fields, for the training, validation, testing, and scaling datasets are shown in Table˜9.
@@
Table 9: Performance metrics for predicting with different models and datasets. Errors in [°C] or [m/y], MoC in [%], and SSIM unitless.
Architecture Data Split Scaling Step MAE RMSE Huber SSIM MoC
UNet synthetic: 3dp train No end-to-end 0.1181 0.5537 0.0558 0.8454 2.29
UNet synthetic: 3dp val No end-to-end 0.1930 0.5795 0.0742 0.6515 11.32
UNet synthetic: 3dp test No end-to-end 0.1908 0.6190 0.0736 0.6671 10.85
DilatedCNN synthetic: 3dp train No end-to-end 0.1180 0.2494 0.0270 0.5483 11.28
DilatedCNN synthetic: 3dp val No end-to-end 0.1273 0.2513 0.0287 0.5183 10.67
DilatedCNN synthetic: 3dp test No end-to-end 0.1200 0.2236 0.0237 0.5004 10.95
PINN (overfit) synthetic: 1dp train No 3 0.7532 0.8593 0.3454 0.1793 0
LGCNN synthetic: 3dp train No end-to-end 0.0657 0.1640 0.0121 0.8209 0.08
LGCNN synthetic: 3dp val No end-to-end 0.0965 0.1996 0.0184 0.6933 1.54
LGCNN synthetic: 3dp test No end-to-end 0.0916 0.1738 0.0146 0.6841 0.09
LGCNN synthetic: 3dp test Yes end-to-end 0.0413 0.1189 0.0065 0.7911 0.06
UNet synthetic: 3dp train No 1 (vx) 9.9620 13.0768 9.4732 0.9868 -
UNet synthetic: 3dp val No 1 (vx) 22.7179 33.2020 22.2241 0.9748 -
UNet synthetic: 3dp test No 1 (vx) 22.3178 31.1860 21.8237 0.9739 -
UNet synthetic: 3dp test Yes 1 (vx) 24.9261 34.7004 24.4314 0.9784 -
UNet synthetic: 3dp train No 1 (vy) 11.8902 16.6018 11.4005 0.9890 -
UNet synthetic: 3dp val No 1 (vy) 27.1026 39.0450 26.6078 0.9770 -
UNet synthetic: 3dp test No 1 (vy) 32.7444 45.0703 32.2488 0.9733 -
UNet synthetic: 3dp test Yes 1 (vy) 26.2795 38.2599 25.7847 0.9761 -
UNet synthetic: 3dp train No 3 0.0282 0.0632 0.0020 0.9376 0.11
UNet synthetic: 3dp val No 3 0.0452 0.0869 0.0038 0.8454 0.11
UNet synthetic: 3dp test No 3 0.0417 0.0762 0.0029 0.8540 0.07
UNet synthetic: 3dp test Yes 3 0.0168 0.0373 0.0007 0.8895 0.08
UNet, in sequence synthetic: 3dp train No 3 0.0402 0.1039 0.0052 0.9019 0.08
UNet, in sequence synthetic: 3dp val No 3 0.1013 0.2025 0.0191 0.6682 1.15
UNet, in sequence synthetic: 3dp test No 3 0.0919 0.1695 0.0139 0.6714 1.30
UNet, with zero-padding synthetic: 3dp train No 3 0.0440 0.0788 0.0031 0.8104 0.44
UNet, with zero-padding synthetic: 3dp val No 3 0.0548 0.0908 0.0041 0.7546 0.70
UNet, with zero-padding synthetic: 3dp test No 3 0.0487 0.0795 0.0031 0.7784 0.57
UNet, no partitioning synthetic: 3dp train No 1 (vx) 3.6714 4.6374 3.1983 0.9959 -
UNet, no partitioning synthetic: 3dp val No 1 (vx) 25.7367 37.7048 25.2418 0.9764 -
UNet, no partitioning synthetic: 3dp test No 1 (vx) 23.8584 33.8297 23.3646 0.9765 -
UNet, no partitioning synthetic: 3dp train No 1 (vy) 3.5177 4.6358 3.0496 0.9971 -
UNet, no partitioning synthetic: 3dp val No 1 (vy) 30.1247 42.7471 29.6294 0.9762 -
UNet, no partitioning synthetic: 3dp test No 1 (vy) 35.9249 49.9040 35.4289 0.9721 -
UNet, no partitioning synthetic: 3dp train No 3 0.0323 0.0736 0.0027 0.9245 0.14
UNet, no partitioning synthetic: 3dp val No 3 0.0449 0.0861 0.0037 0.8333 0.12
UNet, no partitioning synthetic: 3dp test No 3 0.0429 0.0793 0.0031 0.8349 0.29
UNet real: 4dp train No 1 (vx) 14.1819 18.4573 13.6890 0.9915 -
UNet real: 4dp val No 1 (vx) 15.4095 19.5032 14.9187 0.9869 -
UNet real: 4dp test Yes 1 (vx) 110.0078 116.4903 109.5079 0.9436 -
UNet real: 4dp train No 1 (vy) 9.5619 11.2571 9.0680 0.9945 -
UNet real: 4dp val No 1 (vy) 10.6605 12.1757 10.1675 0.9932 -
UNet real: 4dp test Yes 1 (vy) 17.6051 24.8298 17.1118 0.9905 -
UNet real: 4dp train No 3 0.0139 0.0212 0.0002 0.9361 1.88
UNet real: 4dp val No 3 0.0175 0.0319 0.0005 0.9405 3.30
UNet real: 4dp test Yes 3 0.0189 0.0287 0.0004 0.8942 1.77
LGCNN real: 4dp train No end-to-end 0.0534 0.0997 0.0049 0.6900 2.76
LGCNN real: 4dp val No end-to-end 0.0841 0.1659 0.0137 0.6511 1.65
LGCNN real: 4dp test Yes end-to-end 0.0394 0.0665 0.0022 0.6755 3.50
@@
The model generally obtains good results in Figure˜3, even for cells that are far away from injection points.
@@
Additional tests show that excluding other inputs, either $i$ alone or both $i$ and $k$, raises prediction error by 58–121%. We also evaluated alternative time-stepping schemes for solving the IVP. Replacing the 5th-order implicit Runge–Kutta method with explicit 2nd- or 4th-order schemes accelerates computation, but increases prediction error by 16–23%—a moderate degradation compared to the complete removal of streamline inputs. Nonetheless, we retain the implicit scheme for its superior accuracy and stability. Quantitative and qualitative results for the predictions are shown in Figure˜7 and Table˜10.
Table 10: Experiment on 3rd step: Test metrics for predicting $T$ with different input combinations. Errors in [°C], MSE in [°C2], and SSIM unitless.
Inputs Huber L MAE MSE SSIM
$ikv_{x}v_{y}$ 0.0070 2.2990 0.0712 0.0139 0.7662
$ikv_{x}v_{y}s$ 0.0041 1.8674 0.0545 0.0083 0.8368
$ikv_{x}v_{y}s_{o}$ 0.0057 2.6623 0.0598 0.0114 0.8423
$ikv_{x}v_{y}ss_{o}$ (not faded) 0.0072 2.3039 0.0744 0.0144 0.7837
$ikv_{x}v_{y}ss_{o}^{\mathrm{a}}$ 0.0031 1.8364 0.0442 0.0062 0.8828
$v_{x}v_{y}ss_{o}$ 0.0066 2.1301 0.0647 0.0132 0.8681
$kv_{x}v_{y}ss_{o}$ 0.0049 2.0925 0.0587 0.0097 0.8732
explicit RK, order 4 0.0038 3.2636 0.0486 0.0076 0.8871
explicit RK, order 2 0.0036 2.2034 0.0463 0.0072 0.8830
@@
with Huber validation loss, hence the results differ slightly wrt. to Table˜1.
Figure 7: 1st Column: Label, input $k$, 2nd-5th: Without $(s,s_{o})$, $(s_{o})$, $(s)$, not-faded streamlines $(s,s_{o})$, 6th: include all inputs, 7th-8th: Without $(i,k)$, $(i)$. Absolute errors capped at 1°C for better visualizations. Maximum errors are listed in Table˜10.
@@
The qualitative performance is observable in Figure˜8, where we see coherent streamlines and plume structures for both the isolated 3rd step and the full pipeline.
@@
Table˜11 summarizes the training and inference times, number of epochs, and dataset splits (train:val:test) for each of the three steps in our pipeline, both for the LGCNN trained on partitioned and full datasets. Although both approaches exhibit similar inference times, they show significant differences in training time, with the partitioned approach yielding better results.
Table 11: Training measurements on 3dp dataset.
@@
In Table˜12, the number of epochs and the total training time for the data-driven approaches are shown. For Step 1, both the UNet and DDUNet need many epochs and comparable training time to converge; however, for the third step and the full pipeline, the DDUNet significantly reduces both the number of epochs and the total training time required to reach convergence.
@@
$\times2$ DDUNet101dp 735 4.787 hours
@@
$\times2$ DDUNet101dp 303 3.508 hours
@@
$\times2$ DDUNet101dp 97 1.222 hours